# Cutting is All You Need: Execution of Large-Scale Quantum Neural Networks on Limited-Qubit Devices

### v1

@@ Abstract
The rapid advancement in Quantum Computing (QC), particularly through Noisy-Intermediate Scale Quantum (NISQ) devices, has spurred significant interest in Quantum Machine Learning (QML) applications. Despite their potential, fully-quantum QML algorithms remain impractical due to the limitations of current NISQ devices. Hybrid quantum-classical neural networks (HQNNs) have emerged as a viable alternative, leveraging both quantum and classical computations to enhance machine learning capabilities. However, the constrained resources of NISQ devices, particularly the limited number of qubits, pose significant challenges for executing large-scale quantum circuits.
This work addresses these current challenges by proposing a novel and practical methodology for quantum circuit cutting of HQNNs, allowing large quantum circuits to be executed on limited-qubit NISQ devices. Our approach not only preserves the accuracy of the original circuits but also supports the training of quantum parameters across all subcircuits, which is crucial for the learning process in HQNNs. We propose a cutting methodology for HQNNs that employs a greedy algorithm for identifying efficient cutting points, and the implementation of trainable subcircuits, all designed to maximize the utility of NISQ devices in HQNNs. The findings suggest that quantum circuit cutting is a promising technique for advancing QML on current quantum hardware, since the cut circuit achieves comparable accuracy and much lower qubit requirements than the original circuit.
@@ I Introduction
Recent advances in Quantum Computing (QC), particularly with Noisy-Intermediate Scale Quantum (NISQ) devices, have shown groundbreaking progress and demonstrated quantum advantages in various fields [1, 2, 3, 4, 5, 6]. However, due to the limitations of NISQ devices, standalone fully-quantum algorithms remain impractical. Hybrid Quantum-Classical Neural Networks (HQNNs) have emerged as an effective alternative, combining classical and quantum layers for machine learning applications [7, 8, 9, 10]. HQNNs are robust against errors in NISQ devices [11, 12] and employ quantum circuits that interface between classical and quantum computations [13]. Angle embedding and amplitude embedding are commonly used to map classical data into quantum states [14, 15]. The angle embedding method consists of rotating each classical feature by a certain angle based on its magnitude, while the amplitude embedding encodes classical data into the amplitudes of a quantum state. The angle embedding, due to its 1-to-1 representation with the number of features into qubit states, is known for its high precision. On the other hand, using the amplitude encoding, $n$ qubits are sufficient to encode $2^{n}$ features [14]. Hence, it can encode the information in a compact form.
@@ I-A Target Research Problem and Associated Challenges
The limited number of qubits and noisy nature of NISQ devices pose challenges in scaling quantum circuits [16]. Large-scale quantum circuits often suffer from noise, Barren plateaus [17], and computational overhead. The most prominent solutions for mitigating the noise in quantum circuits, namely quantum error correction and mitigation, further increase the demand for many-qubit circuits, since they introduce redundancies in quantum systems to reduce/eliminate the noise [18, 19].
Hence, there is a dire need to execute large-scale quantum circuits onto limited-qubit NISQ devices. Even the classical simulation of quantum circuits becomes extremely computationally and memory-intensive as well as requires very long execution times when simulating large-scale quantum systems. In this regard, Quantum circuit cutting [20] is a promising technique that partitions large circuits into smaller subcircuits, making it feasible to execute them on limited-qubit devices. However, existing quantum circuit cutting techniques [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34] do not support gradient-based training for HQNNs [^1] The current implementation of the automatic differentiation of quantum circuits [13] supports the conversion of a single quantum node (QNode) into Torch/Keras layer, but the cut circuit cannot be mapped onto a single QNode because it requires multiple nodes.. Our work introduces a methodology that supports quantum circuit cutting while maintaining the backward propagation of gradients across subcircuits.
I-B Motivational Analysis
@@
As a motivational case study, we analyze the training process of four different variants of HQNN models. We assume that the HQNN model in Figure 1a has the best architecture for the given application, which in this analysis is the classification on the Digits dataset [35]. This architecture is composed of a set of classical layers and quantum layers. A classical fully-connected (FC) layer with 8 neurons is followed by an 8-qubit quantum circuit and a classical FC layer with 10 neurons. The 8 classical features at the output of the first classical FC layer are encoded into quantum states with angle embedding. The final quantum states are then measured, and their values are used as input features of the following classical FC layer. The implementation, training, and hyperparameter details are discussed in Section IV-A.
However, we assume that only a 3-qubit quantum device is available. In this regard, Figure 1b, c, and d present three possible alternatives [^2] These three solutions do not form an exhaustive list of possibilities, but they provide a variety of solutions to solve the problem. to implement a modified version of the original circuit in a 3-qubit device. The 3-3 Angle architecture (version b) reduces the size of the first classical FC layer to 3 and maintains angle embedding to encode the three features into 3-qubit states. The 8-3 Amplitude architecture (version c) employs amplitude embedding to encode the 8 classical features into 3 quantum states in the 3-qubit device. The 8-3 Cut architecture (version d) employs the quantum circuit cutting technique to implement the 8-qubit circuit into multiple 3-qubit subcircuits in the 3-qubit quantum device.
The results of training and validation accuracy shown in Figure 2 highlight that the 8-3 Cut architecture provides the best solution in terms of accuracy compared to 3-3 Angle and 8-3 Amplitude. As indicated by labels  1 and  3, the 8-3 Cut HQNN learns faster than the original 8-8 Angle HQNN in the first few training epochs. On the other hand, both 3-3 Angle and 8-3 Amplitude architectures do not reach the same level of accuracy of the original 8-8 Angle HAQNN (see label  2). The lower accuracy of the 3-3 Angle HQNN is due to an information loss because only 3 features are propagated after the first FC layer, instead of 8. The 8-3 Amplitude HQNN has lower accuracy than the original 8-8 Angle and 8-3 Cut because of the different type of embedding used to encode the data. Moreover, as highlighted by label  4, the 8-3 Amplitude HQNN requires a larger number of epochs than other architectures to reach high validation accuracy.
Figure 2: (a) Training and (b) validation accuracy for different variants of the HQNN architecture on the Digits dataset.
This simple yet insightful analysis demonstrates that the approach with quantum circuit cutting has the potential to execute quantum circuits on limited-qubit devices with minimal changes to the rest of the HQNN architecture, maintaining similar accuracy levels compared to the original circuit.
@@ I-C Novel Contributions
  • We propose a methodology for cutting HQNNs into multiple subcircuits, to enable their full execution on limited-qubit devices. (Section III)
  • We design a greedy algorithm to automatically identify the cutting points in a given quantum circuit under a given qubit constraint for the underlying quantum device. (Section III-A)
@@
  • We evaluate our methodology on various benchmarks and configurations of quantum circuits and qubit requirements. The experimental results demonstrate that the cut circuits achieve comparable accuracy to the original circuit with a small computation overhead, thereby enabling the execution of more complex circuits on limited-qubit devices. (Section IV)
Figure 3: Overview of the novel contributions in this work.
@@ II-A Hybrid Quantum-Classical Neural Networks
HQNNs are hybrid systems that combine classical and quantum computational elements, making them well-suited for NISQ devices [36, 37, 38]. In an HQNN, classical and quantum layers are integrated, typically in a sequential manner, allowing for a synergistic use of both types of computational resources [39]. Classical layers handle preprocessing, feature extraction, and output, while quantum layers perform complex operations like quantum feature mapping and entanglement. Similarly to the weights of the classical layers, the gates of the parameterized quantum circuits (PQCs) contain trainable parameters [40].
Classical data is transitioned into quantum layers by encoding it into quantum states through techniques such as angle embedding, which maps features to qubit rotations, or amplitude embedding, which compresses data into the quantum state’s amplitudes [14]. The quantum layers then process this data, leveraging high-dimensional feature spaces to analyze complex patterns. After quantum operations, the results are measured and fed back to the classical layers for further processing or decision-making [41]. The hybrid structure of HQNNs takes advantage of quantum computing’s strengths, such as parallelism and entanglement, while the classical layers mitigate the limitations of current quantum hardware [42]. This design enhances robustness against noise and errors and allows for greater scalability, making HQNNs a promising approach for advancing quantum machine learning, especially when fully-quantum networks are not yet feasible [36].
@@ II-B Quantum Circuit Cutting
In recent years, several variants and optimizations of the quantum circuit cutting method have been proposed in the literature [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. Moreover, quantum circuit cutting has been used for error mitigation [44, 45, 46].
@@
On the other hand, the works in [47, 48] are applied to QML but they only focus on quantum circuit partitioning. Similarly, the works in [49, 50] describe methods to execute HQNNs on distributed systems. However, the scope of our work is to provide a platform for employing quantum circuit cutting on HQNNs.
@@ 1 Procedure CuttingPointsDesign(Gates, n)
2       Sort Gates in Priority Order, from $1$ to $MaxGate$; // priority based on the data dependencies in the circuit
@@ IV-B Evaluation of our Cutting Methodology on the Digits Dataset
The results of cutting the circuits for the Digits dataset are shown in Figure 8. The results contain a set of $n$-$m$ cut experiments, comparing the training and validation accuracy of the original $n$-qubit circuit with the cut circuit implemented on an $m$-qubit device. For different values of $m$ and $n$, different observations can be derived. As indicated in labels  12, and  3, for $4$-$2$, $6$-$2$, and $8$-$2$ cuts there is quite a large gap between the training accuracy of the original circuit and the training accuracy of the cut circuit. A similar large gap can be observed for the validation accuracy. The reason for this discrepancy can be attributed to the fact that when the circuits are implemented on $2$ qubits, the $2$-qubit subcircuits have low correlation and entanglement capabilities, which hinder the correct generalization of the problem.
However, for the $10$-$2$ cut circuit, the results are slightly different (see label  4). While the training and validation accuracy gaps between the original circuit and the cut circuit are quite large after a few epochs, they shrink towards the end of the training epochs. This behavior suggested that the training process of the cut citcuits implemented on a $2$-qubit device might require more training epochs to reach higher accuracy. Hence, in Section IV-E, we analyze the behavior of these cut circuits when the HQNNs are trained for more epochs.
On the other hand, for devices that have more than $2$ qubits, the training and validation accuracy of the cut circuits are comparable with the original circuit. As highlighted in labels  5 and  6 in Figure 8, for the $8$-$3$ cut and $10$-$3$ cut circuits, during the first few epochs, the training and validation accuracy of the cut circuits are higher than the original circuits, meaning that the cut circuits converge faster than the original circuit. Moreover, the validation accuracy of the cut circuit for the $10$-$3$ cut even results slightly higher than the validation accuracy of the original circuits (see label  7). This can be attributed to a better generalization when the cuts are deployed in the circuit. For other cases (see labels  8 and  9), the training and validation accuracy curves for the original and cut circuits are overlapping, indicating that there is no significant functional difference between the two implementations.
@@ IV-C Evaluation of our Cutting Methodology on the MNIST Dataset
Figure 9 shows the comparative results between the cut circuits and original circuits for the MNIST dataset. Overall, we can observe that the training and validation accuracy curves are similar for all the configurations. In some cases, for instance, for the $8$-$4$ cut (see label  1 in Figure 9), the training and validation accuracy of the cut circuit is slightly lower than that of the original circuit. In other cases, for the $6$-$4$ cut and $10$-$3$ cut circuits (see labels  2 and  3). Moreover, we can observe that for the first few epochs, the training and validation accuracies for the $8$-$3$ cut circuit are higher than for the original circuit (see label  4).
@@ IV-D Evaluation of the Impact of FLOPs Variations
IV-E Ablation Study: Training Cut Circuits for More Epochs
Since the slope of the learning curves in some of the cut circuits is still high after 50 training epochs, we extended the training process to 200 epochs. Figure 10 shows the training and validation accuracy curves for the $4$-$2$, $6$-$2$, $8$-$2$, and $10$-$2$ cut circuits. We can observe that the accuracy plateau is reached after 120 epochs for the $6$-$2$ and $8$-$2$ circuits (see labels  1 and  2 in Figure 10). For higher qubits of the original circuit (i.e., for the $10$-$2$ cut), the convergence appears after fewer epochs, while for lower qubits of the original circuits (i.e., for the $4$-$2$ cut), more training epochs are required to obtain the convergence.
Figure 10: (a) Traning and (b) validation accuracy of 4-2, 6-2, 8-2, and 10-2 cut circuits, trained for 200 epochs on the Digits dataset.
IV-F Discussion

### v2

@@ Abstract
The rapid advancement in Quantum Computing, particularly through Noisy-Intermediate Scale Quantum (NISQ) devices, has spurred significant interest in Quantum Machine Learning (QML) applications. Despite their potential, fully-quantum algorithms remain impractical due to the limitations of current NISQ devices. Hybrid quantum-classical neural networks (HQNNs) have emerged as a viable alternative, leveraging both quantum and classical computations to enhance machine learning capabilities. However, the constrained resources of NISQ devices, particularly the limited number of qubits, pose significant challenges for executing large-scale quantum circuits.
This work addresses these current challenges by proposing a novel and practical methodology for quantum circuit cutting of HQNNs, allowing large quantum circuits to be executed on limited-qubit NISQ devices. Our approach not only preserves the accuracy of the original circuits but also supports the training of quantum parameters across all subcircuits, which is crucial for the learning process in HQNNs. We propose a cutting methodology for HQNNs that employs a greedy algorithm for identifying efficient cutting points, and the implementation of trainable subcircuits, all designed to maximize the utility of NISQ devices in HQNNs. The findings suggest that quantum circuit cutting is a promising technique for advancing QML on current quantum hardware, since the cut circuit achieves comparable accuracy and much lower qubit requirements than the original circuit. The code is available at https://github.com/eBrain4Everyone/QNN-Cutting.
Index Terms:
Quantum Neural Networks, Circuit Cutting, Quantum Machine Learning
@@ I Introduction
Quantum computing has demonstrated significant advancements, particularly with Noisy-Intermediate Scale Quantum (NISQ) devices [1, 2, 3, 4, 5, 6, 7, 8]. However, due to the constraints of NISQ hardware, such as limited qubits, noise, and scalability issues, fully Quantum Machine Learning (QML) algorithms remain impractical [9, 10]. Hybrid Quantum-Classical Neural Networks (HQNNs) have emerged as a promising alternative, combining classical and quantum layers for machine learning tasks [11, 12, 13, 14]. HQNNs leverage quantum circuits for complex representations while retaining classical robustness, making them suitable for NISQ-era hardware [15, 16].
@@ I-A Target Research Problem and Associated Challenges
Despite their potential, HQNNs face scalability challenges due to qubit limitations. Large-scale quantum circuits suffer from noise[17, 18, 19, 20], barren plateaus [21, 22, 23, 24, 25], and significant computational overhead. Quantum circuit cutting [26] is a promising technique that partitions large circuits into smaller subcircuits, enabling execution on limited-qubit devices.
However, while circuit cutting has shown success in other quantum computing applications, applying it to trainable models like HQNNs introduces a fundamentally different set of challenges. Existing cutting techniques [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40] are incompatible with gradient-based training because they lack support for end-to-end differentiability. This limitation is critical: training HQNNs requires propagating gradients through quantum subcircuits, a feature incompatible with traditional cutting approaches. For example, current automatic differentiation frameworks like PennyLane [41] are limited to differentiating a single quantum node (QNode) by converting it into a differentiable Torch or Keras layer. However, a cut circuit in HQNNs spans multiple QNodes, breaking the gradient flow required for backpropagation.
This gap motivates the need for a separate and focused investigation into circuit cutting tailored for HQNNs. Unlike conventional circuit cutting, which optimizes for execution or fidelity, cutting in the context of HQNNs must preserve differentiability, model integrity, and training dynamics. As such, circuit cutting for HQNNs is not merely an extension of prior techniques, but it is a distinct problem that necessitates new strategies for partitioning and gradient management.
Summary of Challenges:
  • Limited qubit availability: Large-scale quantum circuits require more qubits than available in current NISQ devices.
  • Gradient propagation in cut circuits: Existing quantum circuit cutting techniques do not support backpropagation across subcircuits, preventing effective training of HQNNs.
  • Computational efficiency: While circuit cutting enables execution on limited-qubit devices, it introduces overhead, requiring an optimized approach for balancing accuracy and efficiency.
To overcome these challenges, we propose a framework that (i) enables quantum circuit cutting for HQNNs, (ii) maintains trainability by supporting gradient backpropagation across subcircuits, and (iii) efficiently partitions circuits under hardware constraints.
I-B Motivational Case Study
@@
We evaluate four HQNN variants on the Digits dataset [42] (Fig. 1). The original HQNN (8-8 Angle) contains an 8-neuron fully connected (FC) layer, followed by an 8-qubit quantum circuit (QCt), and a 4-neuron FC layer. The 8 classical features at the output of the first classical FC layer are encoded into quantum states with angle embedding [^1] Angle embedding and amplitude embedding are commonly used to map classical data into quantum states [43, 44]. The angle embedding method consists of rotating each classical feature by a certain angle based on its magnitude, while the amplitude embedding encodes classical data into the amplitudes of a quantum state. The angle embedding, due to its 1-to-1 representation with the number of features into qubit states, is known for its high precision. On the other hand, using the amplitude encoding, $n$ qubits are sufficient to encode $2^{n}$ features [43]. Hence, it can encode the information in a compact form.. Given an available 3-qubit quantum device, we explore:
  • 3-3 Angle: Reducing the first FC layer to 3 neurons, maintaining angle embedding.
  • 8-3 Amplitude: Encoding 8 features into 3 qubits using amplitude embedding.
  • 8-3 Cut: Applying quantum circuit cutting to execute an 8-qubit circuit on a 3-qubit device.
The experimental results shown in Figure 2 demonstrate that the 8-3 Cut architecture preserves accuracy while reducing qubit requirements. The 3-3 Angle and 8-3 Amplitude architectures suffer from feature loss and encoding inefficiencies, leading to lower accuracy (see label  2). As indicated by labels  1 and  3, the 8-3 Cut HQNN converges faster than the original 8-8 Angle HQNN in early training epochs. The lower accuracy of the 3-3 Angle HQNN stems from information loss, as only 3 features are propagated instead of 8. The 8-3 Amplitude HQNN underperforms compared to the 8-8 Angle and 8-3 Cut architectures due to the different embedding strategy. Moreover, as highlighted by label  4, it requires more epochs to achieve high validation accuracy.
Figure 2: Training and validation accuracy for different HQNN architectures.
This analysis highlights that quantum circuit cutting enables the execution of quantum circuits on limited-qubit devices while maintaining accuracy comparable to the original HQNN architecture.
@@ I-C Novel Contributions
Figure 3: Overview of the novel contributions in this work.
  • We propose a framework for cutting HQNNs into multiple subcircuits, to enable their full execution on limited-qubit devices. (Section III)
  • We design a greedy algorithm to identify the cutting points based on the given qubit constraints. (Section III-A)
@@
  • We evaluate our methodology across various benchmarks, demonstrating accuracy preservation with reduced qubit requirements. (Section IV)
  • For reproducible research, we release a repository containing the code of the complete framework with cut circuits at https://github.com/eBrain4Everyone/QNN-Cutting.
By enabling large-scale HQNNs on constrained quantum hardware, our approach paves the way for practical QML in the NISQ era.
@@ II-A Hybrid Quantum-Classical Neural Networks
HQNNs are hybrid systems that combine classical and quantum computational elements, making them well-suited for NISQ devices [45, 46, 47]. In an HQNN, classical and quantum layers are integrated, typically in a sequential manner, allowing for a synergistic use of both types of computational resources [48, 49, 50]. Classical layers handle preprocessing, feature extraction, and output, while quantum layers perform complex operations like quantum feature mapping and entanglement. Similarly to the weights of the classical layers, the gates of the parameterized quantum circuits (PQCs) contain trainable parameters [51].
Classical data is transitioned into quantum layers by encoding it into quantum states through techniques such as angle embedding, which maps features to qubit rotations, or amplitude embedding, which compresses data into the quantum state’s amplitudes [43]. The quantum layers then process this data, leveraging high-dimensional feature spaces to analyze complex patterns. After quantum operations, the results are measured and fed back to the classical layers for further processing or decision-making [52]. The hybrid structure of HQNNs takes advantage of quantum computing’s strengths, such as parallelism and entanglement, while the classical layers mitigate the limitations of current quantum hardware [53, 54]. This design enhances robustness against noise and errors and allows for greater scalability, making HQNNs a promising approach for advancing quantum machine learning, especially when fully-quantum networks are not yet feasible [45].
@@ II-B Quantum Circuit Cutting
In recent years, several variants and optimizations of the quantum circuit cutting method have been proposed in the literature [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. Moreover, quantum circuit cutting has been used for error mitigation [56, 57, 58] and intellectual property protection [59].
@@
On the other hand, the works in [60, 61] are applied to QML but they only focus on quantum circuit partitioning. The works in [62, 63, 64] discuss how to execute HQNNs on distributed systems. However, the scope of our work is to provide a platform for employing quantum circuit cutting on HQNNs.
@@ 1 Procedure CuttingPointsDesign(Gates, n)
2  Sort Gates in Priority Order, from 11 to MaxGateMaxGate; // priority based on the data dependencies in the circuit
@@ IV-B Evaluation of our Cutting Methodology on the Digits Dataset
The results of cutting the circuits for the Digits dataset are shown in Figure 8. The results contain a set of $n$-$m$ cut experiments, comparing the training and validation accuracy of the original $n$-qubit circuit with the cut circuit implemented on an $m$-qubit device. For different values of $m$ and $n$, different observations can be derived. As indicated in labels  12, and  3, for 44-22, 66-22, and 88-22 cuts there is quite a large gap between the training accuracy of the original circuit and the training accuracy of the cut circuit. A similar large gap can be observed for the validation accuracy. The reason for this discrepancy can be attributed to the fact that when the circuits are implemented on $2$ qubits, the $2$-qubit subcircuits have low correlation and entanglement capabilities, which hinder the correct generalization of the problem.
However, for the 1010-22 cut circuit, the results are slightly different (see label  4). While the training and validation accuracy gaps between the original circuit and the cut circuit are quite large after a few epochs, they shrink towards the end of the training epochs. This behavior suggested that the training process of the cut circuits implemented on a $2$-qubit device might require more training epochs to reach higher accuracy. Hence, in Section IV-F, we analyze the behavior of these cut circuits when the HQNNs are trained for more epochs.
On the other hand, for devices that have more than $2$ qubits, the training and validation accuracy of the cut circuits are comparable with the original circuit. As highlighted in labels  5 and  6 in Figure 8, for the 88-33 cut and 1010-33 cut circuits, during the first few epochs, the training and validation accuracy of the cut circuits are higher than the original circuits, meaning that the cut circuits converge faster than the original circuit. Moreover, the validation accuracy of the cut circuit for the 1010-33 cut even results slightly higher than the validation accuracy of the original circuits (see label  7). This can be attributed to a better generalization when the cuts are deployed in the circuit. For other cases (see labels  8 and  9), the training and validation accuracy curves for the original and cut circuits are overlapping, indicating that there is no significant functional difference between the two implementations.
@@ IV-C Evaluation of our Cutting Methodology on the MNIST Dataset
Figure 9 shows the comparative results between the cut circuits and original circuits for the MNIST dataset. Overall, we can observe that the training and validation accuracy curves are similar for all the configurations. In some cases, for instance, for the 88-44 cut (see label  1 in Figure 9), the training and validation accuracy of the cut circuit is slightly lower than that of the original circuit. In other cases, for the 66-44 cut and 1010-33 cut circuits (see labels  2 and  3), the training and validation accuracy of the cut circuit is slightly higher than that of the original circuit. Moreover, we can observe that for the first few epochs, the training and validation accuracies for the 88-33 cut circuit are higher than for the original circuit (see label  4).
@@ IV-D Evaluation of the Impact of FLOPs Variations
IV-E Evaluation of 20-5 Cut & 50-5 Cut on the MNIST Dataset
Figure 10 presents the training and validation accuracy curves on the MNIST dataset for cut circuits derived from original circuits with 20 and 50 qubits [^4] Only results from the cut circuit are reported for the 50-qubit case.. All cut circuits are implemented using 4 qubits and Angle Embedding. Despite some fluctuations during training (see labels  2 and  4), the curves demonstrate that the cutting technique remains effective even for larger-scale circuits, as accuracy converges to a high and stable level (see labels  1 and  3). Notably, the 20-4 cut circuit shows a significant improvement in both accuracy and convergence speed compared to its uncut 20-qubit counterpart (see labels  1 and  2), highlighting the advantages of circuit cutting.
Figure 10: Training and validation accuracy results for 20 qubits, 20-4 cut, and 50-5 cut on the MNIST dataset.
IV-F Ablation Study: Training Cut Circuits for More Epochs
Since the slope of the learning curves in some of the cut circuits is still high after 50 training epochs, we extended the training process to 200 epochs. Figure 11 shows the training and validation accuracy curves for the 44-22, 66-22, 88-22, and 1010-22 cut circuits. We can observe that the accuracy plateau is reached after 120 epochs for the 66-22 and 88-22 circuits (see labels  1 and  2 in Figure 11). For higher qubits of the original circuit (i.e., for the 1010-22 cut), the convergence appears after fewer epochs, while for lower qubits of the original circuits (i.e., for the 44-22 cut), more training epochs are required to obtain the convergence.
Figure 11: Training and validation accuracy of 4-2, 6-2, 8-2, and 10-2 cut circuits, trained for 200 epochs on the Digits dataset.
IV-G Ablation Study: Amplitude Embedding vs. Angle Embedding
Figure 12 presents a comparative analysis of encoding methods for mapping classical features into quantum states. The results indicate that Angle Embedding outperforms Amplitude Embedding, yielding higher learning curves. Specifically, the training and validation accuracy of the cut circuit with Angle Embedding closely match those of the original circuit (see label  1 in Figure 12). In contrast, Amplitude Embedding results in a noticeable accuracy decline even in the original circuit (see label  2), with further degradation introduced by circuit cutting (see labels  3 and  4). These findings underscore the superiority of Angle Embedding as a more effective encoding strategy for HQNN circuit cutting.
Figure 12: Comparison between Amplitude Embedding and Angle Embedding for 8-3 cuts on the Digits dataset.
IV-H Discussion